# Autoregressive Semantic Visual Reconstruction Helps VLMs Understand Better

### v1

@@ 5.5 Ablation Study
Table 4: Ablation study for various ASVR configurations. This table presents a comparison of various ASVR settings, including semantic tokenizer, varied the degree of alignment with text (e.g., DualToken-12M vs. DualToken-3M [43]), and the training strategy, where "PT/IT" denotes that semantic visual supervision is applied during both the pre-training and instruction tuning stages, while "IT" indicates that semantic visual supervision is applied only during instruction tuning.
@@
Semantic Tokenizer DualToken-12M DualToken-3M 57.8(-1.7) 25.4(+1.1) 33.1(-2.3) 16.2(-0.2) 67.2(+1.1) 70.3(-2.5) 64.8(-1.6) 60.0(-1.5) 31.8(-2.1) 55.9(-1.1) 54.3(+0.2) 24.7(-5.3) 33.0(-0.7) 86.1(-0.2) 48.6
@@
Increasing the degree of alignment with text for semantic tokenizer leads to performance of ASVR. we use different semantic tokenizers to construct semantic visual supervision targets: DualToken-3M, which achieves zero-shot ImageNet classification accuracy of 78.6% [7], and DualToken-12M, which achieves 81.6% and thus exhibits stronger semantic alignment. As shown in Table 4, ASVR equipped with the better-aligned DualToken-12M consistently outperforms the variant using DualToken-3M across the majority of multimodal benchmarks, with the average performance improving by more than 2%. These results demonstrate that employing better semantically aligned visual tokenizer provides semantic visual supervision targets with more meaningful aspects of the image, and further support our claim that Semantic Visual Reconstruction plays a key role in enhancing the multimodal understanding capabilities of LVLMs.
@@
5.6 Qualitative Comparison
Figure 3: Qualitative comparison on attention maps, where we keep the same LLM and training data. With extra vision-centric supervision signals, ROSS urges the model to focus on specific image contents corresponding to the question with higher attention values.
6 Conclusion
In summary, we introduced Autoregressive Semantic Visual Reconstruction (ASVR), enabling joint learning of visual and textual modalities within a unified autoregressive framework and effectively improving multimodal understanding capability of LVLMs. Unlike conventional LVLMs framework, which predominantly rely on textual autoregressive supervision and frequently neglect crucial visual details, ASVR explicitly integrates semantic visual supervision to foster deep perception of visual inputs. Our findings indicate that reconstructing raw visual appearance autoregressively does not benefit, and can even impair multimodal understanding. Conversely, autoregressively reconstructing semantic visual representations of images consistently enhances performance across diverse multimodal tasks. Remarkably, even with continuous visual features as input, ASVR effectively reconstructs discrete semantic tokens, yielding stable and substantial improvements on various multimodal benchmarks. This effectiveness is robust across different visual feature types, LLM backbone capacities, data scales, and high-resolution scenarios, underscoring ASVR’s adaptability and versatility. Future work aims to incorporate image generation capabilities into ASVR, leveraging unified visual autoregressive supervision to seamlessly integrate understanding and generation, thus broadening applicability across diverse downstream tasks.
Appendix A Appendix
A.1 Training Details
Our detailed training settings and hyper-parameters of ASVR are shown in Table 5. We adopt the same training configuration as the baseline LLaVA-1.5 [27] without any additional modifications and find that ASVR is consistently effective under these settings. Notably, SigLIP [63] encodes each $384\times 384$ input image into the sequence of 729 visual features, which exactly matches the sequence length of discrete visual token indices produced by the DualToken-12M [43] visual tokenizer.
Visual Head Since residual quantization introduces a depth-stacked structure of codes at each visual position $p$, we implement our visual heads based on the depth transformer from RQ-VAE [22]. Unlike the original depth transformer—which employs a single head to predict logits across all depths—we follow the design introduced by [43] and use separate classification heads to compute the logits for residuals at each specific depth. Both heads for appearance tokens and semantic tokens share the same structure, comprising three layers of depth transformers, each accompanied by a dedicated classification head for each depth level.
Given the LLM hidden state $h_{p}$ for visual tokens at position $p$, the depth transformer autoregressively predicts D residual tokens ($r_{p1}$, $r_{p2}$, …, $r_{pD}$). For $d>1$, the input to the depth transformer at depth d, denoted as $I_{pd}$, is defined as the sum of the token embeddings of up to depth $d-1$
$I_{pd}=\sum_{d^{\prime}=1}^{d-1}\mathbf{e}(r_{pd^{\prime}}),$ (5)
The initial input at depth 1 is given by $I_{p1}=h_{p}$. This formulation ensures that the depth transformer incrementally refines the predicted feature representation by leveraging previous estimations up to depth $d-1$.
Table 5: Detailed training hyperparameters of ASVR.
Configuration Stage 1 Stage 2
Visual Semantic Tokenizer Dualtoken-12M
Visual encoder siglip-so400m-patch14-384
Projector 2 Linear layers with GeLU
Image resolution 384 x 384
Learning rate 1e-3{projector,visual head} 2e-5{LLM,projector,visual head}
LR schedule Cosine decay
Weight decay 0
Optimizer AdamW
Warmup ratio 0.03
Epoch 1
Global batch size 256 128
Deepspeed Zero2 Zero2
Max token length 4096
A.2 Evaluation Prompts
All prompts used for evaluation benchmarks are released and summarized in Table6 following Cambrian-1 [45].
Table 6: Listing the prompts used in the evaluation of each benchmark.
Benchmark Prompt
TextVQA [42] Answer the question using a single word or phrase.
DocVQA [37] Answer the question using a single word or phrase.
OCRBench [34] Give the short answer directly.
ChartQA [36] Answer the question using a single number or phrase.
MMBench [33] Answer with the option’s letter from the given choices directly.
MME [13] Answer the question using a single word or phrase.
SEED-Image [24] Answer with the option’s letter from the given choices directly.
GQA [18] Answer the question using a single word or phrase.
MMMU [62] Answer with the option’s letter from the given choices directly.
AI2D [21] Answer with the option’s letter from the given choices directly.
RealworldQA [58] Please answer directly with only the letter of the correct option and nothing else.
MMVP [47] Answer with the option’s letter from the given choices directly.
Hallusionbench [15] Answer the question using a single word or phrase.
POPE [26] Answer the question using a single word or phrase.
A.3 The Impact of Backbone Scaling
We further evaluate the generalization capability of ASVR under the larger-scale LLM backbone. Specifically, we extend our experiments to Vicuna-v1.5-13B[65], The training data is LLaVA-558K [28] and LLaVA-1.5-665K [28] for the pre-training stage and the instruction tuning stage respectively, keeping the training conditions identical. As shown in Table7, ASVR consistently outperforms the baseline across a wide range of multimodal benchmarks, demonstrating its effectiveness in scaling with larger LLM backbones.
Table 7: The Generality of ASVR with LLM backbone scaling across multimoda understanding benchmarks. "✗" indicates the use of textual supervision only, while "✓" denotes the inclusion of semantic visual supervision by computing additional $\mathcal{L}_{\mathrm{AR}}^{\text{vision}}$. Visual encoder(SigLIP-ViT-SO400M/14@384) are both utilized for ASVR and baseline. "/" separates the data scale used for pre-training (left) and instruction tuning (right).
@@ A.3 The Impact of Backbone Scaling
TVQA DVQA OCRB CQA MMB MME SEED GQA MMMU AI2D RQA MMVP Hbench POPE
LLaVA Vicuna-v1.5-13B 558K/665k 57.2 22.1 32.4 15.1 67.1 68.9 65.6 60.4 35.6 54.9 54.8 34.0 32.9 86.8 49.1
A.4 Comparison with ROSS
ROSS[50] applies the denoising objective to reconstruct visual tokens, whereas ASVR adopts autoregressive objective to reconstruct semantic visual tokens. Both approaches aim to construct visual supervision for LVLMs to enhance multimodal understanding capabilities.
Table 8: The performance comparison between ASVR and ROSS under identical training configurations across five representative multimodal understanding tasks, each reflecting a distinct capability dimension.
ChartQA MMBench MMMU RealworldQA Hallusionbench AVG
ROSS 16.2 67.7 32.8 53.5 32.7 40.6
ASVR 16.4 66.1 33.9 54.1 33.7 40.8
In TableA.4, we present the performance comparison between ROSS and ASVR under identical training settings across five multimodal benchmarks, each representing different capability dimension. Notably, the training hyperparameters are directly borrowed from ROSS [51]. Specifically, we configure SigLIP-ViT-SO400M/14@384 [63] as the visual encoder and Vicuna-v1.5-7B [65] as the LLM backbone. Both ASVR and ROSS are trained using LLaVA-558K[28] for the pre-training stage, LLaVA-1.5-665K[28] for the instruction tuning stage.

### v2

@@ A.5 Ablation Study
Table 8: Ablation study for various ASVR configurations. This table presents a comparison of various ASVR settings, including semantic tokenizer, varied the degree of alignment with text, and the training strategy, where "PT/IT" denotes that semantic visual supervision is applied during both the pre-training and instruction tuning stages, while "IT" indicates that semantic visual supervision is applied only during instruction tuning.
@@
Semantic Tokenizer 12M 3M 57.8(-1.7) 25.4(+1.1) 33.1(-2.3) 16.2(-0.2) 67.2(+1.1) 70.3(-2.5) 64.8(-1.6) 60.0(-1.5) 31.8(-2.1) 55.9(-1.1) 54.3(+0.2) 24.7(-5.3) 33.0(-0.7) 86.1(-0.2) 48.6
@@
Increasing the degree of alignment with text for semantic tokenizer leads to performance of ASVR. We use VQ-SigLip (song2025dualtoken) trained on different data scales to construct semantic visual supervision targets: with 3M data trained varient, which achieves zero-shot ImageNet classification accuracy of 78.6% (deng2009imagenet), and with 12M data trained varient, which achieves 81.6% and thus exhibits stronger semantic alignment. As shown in Table 8, ASVR equipped with the better-aligned with 12M data trained varient consistently outperforms the variant using 3M data trained across the majority of multimodal benchmarks, with the average performance improving by more than 2%. These results demonstrate that employing better semantically aligned visual tokenizer provides semantic visual supervision targets with more meaningful aspects of the image, and further support our claim that Semantic Visual Reconstruction plays a key role in enhancing the multimodal understanding capabilities of LVLMs. Moreover, when the supervised visual tokenizer provides richer semantic information, ASVR achieves stronger performance. We present the results obtained using discrete SigLIP2 (tschannen2025siglip2multilingualvisionlanguage) as visual tokenizer which contain richer semantic visual information in the Appendix A.5.
@@
Table 9: Extend experiments on LLaVA-Next dataset, LLaVA indicates the baseline (typically LVLM framework), ASVR builds upon the baseline by introducing autoregressive semantic visual supervision. "✗" indicates the use of textual supervision only. Visual encoder(SigLIP-ViT-SO400M/14@384 and SigLIP2-ViT-SO400M/14@384) are both utilized for ASVR and baseline to get different visual input and visual supervision.
@@
Visual Encoder Visual Supervision LLM Backbone Data TVQA DVQA OCRB CQA MMB MME SEED GQA MMMU AI2D RQA MMVP Hbench POPE AVG
LLaVA Siglip-so400m-384 Vicuna-v1.5–7B LLaVA–Next 57.7 40.7 37.9 42.6 67.4 71.5 67.2 61.8 34.3 65.3 54.6 32.8 33.1 86.4 53.8
ASVR Siglip-so400m-384 Semantic Siglip Vicuna-v1.5–7B LLaVA–Next 58.6 40.6 39.7 43.4 67.9 73.0 67.5 62.9 34.2 65.8 55.4 36.8 39.2 85.9 55.1
LLaVA Siglip2-so400m-384 Vicuna-v1.5–7B LLaVA–Next 59.2 41.8 40.5 46.3 66.9 74.0 68.3 62.7 34.7 66.4 56.9 33.3 35.1 86.1 55.2
ASVR Siglip2-so400m-384 Semantic Siglip-2 Vicuna-v1.5–7B LLaVA–Next 61.0 43.7 44.8 49.9 70.2 76.8 69.5 63.4 36.3 67.3 56.7 42.0 35.8 86.8 57.4
The Impact of visual supervision
We further extend our experiments by employing discrete VQ-SigLIP2 (tschannen2025siglip2multilingualvisionlanguage) as visual supervision, which provides richer and stronger semantic information, to verify that enhanced visual-semantic supervision can better scale the effectiveness of ASVR. To ensure fair comparison, we use LLaVA-Next (liu2024llava) as the training dataset under identical conditions, evaluating ASVR against the baseline with SigLIP (zhai2023siglip) as both visual input and supervision, as well as with SigLIP2 (tschannen2025siglip2multilingualvisionlanguage) serving the same roles.
The results shown in Table 9 clearly demonstrate that stronger visual semantic encoders lead to better performance when used for supervision. Specifically, ASVR with SigLIP-2 outperforms the baseline (LLaVA) with SigLIP-2 by an average of +2.2 points across 14 benchmarks. In comparison, ASVR with SigLIP improves over its baseline by +1.3 points. These results indicate that ASVR benefits more from stronger semantic supervision, and that pairing ASVR with more powerful semantic vision supervision further enhances its ability to improve visual understanding.